OpenStudy (anonymous):

in this problem k^2+8k+12=0 im at (k-16)^2 and then square rooted =s square root of -12 and you cant square root a negative so is it prime or is it k= positive and negative square root of -12,+16?

OpenStudy (anonymous):

how did you get to (k-16)^2? looks like (k+6)another factor. and you can take a negative square root you just have to use I but don't worry about that because it doesn't apply since this factors nicely

OpenStudy (anonymous):

i subtarted the 12 from both sides so k^2+8k+ (space)=-12 and then half of 8 is 4 and 4^2 is 16 so i filled the space with 16 k^2+8k+ 16=-12 and then some how turn that to (k-16) = -12

OpenStudy (anonymous):

you dont need to complete the square on this since it factors easily and when you add that 16 you have to do it to both sides.

OpenStudy (anonymous):

okay so k^2+8k+16=4

OpenStudy (anonymous):

you don't have to complete the square. but if you must. factor the left side, and solve the equation. you will get (k+4)^2=4 k+4=+-2 then finish from there. that means plus or minus by the way. if we decide to take the root we have to consider both possibilities

OpenStudy (anonymous):

hold up can you explain how you get (k=4)^2 ? is it because 4 goes into booth 8 and 16 ?

OpenStudy (anonymous):

its because 4*4=16 and 4+4=8 the same reason we could have just factored this easily to begin with as (k+6)(k+2) 6+2=8 and 6*2=12

OpenStudy (anonymous):

okay ... so k=+-2-4 ?

OpenStudy (anonymous):

and to expand on what you had in the beginning the sqrt(-12)= \[\sqrt{4}*\sqrt{3}i=2\sqrt{3}i\] and yes. that is correct

OpenStudy (anonymous):

almighty thanks again hahah inventively i will catch on and be able to due these on my own ... hopefully before my test Wednesday

OpenStudy (anonymous):

alrighty*

OpenStudy (anonymous):

just do 1000 of them between now and then and you should be good. your best bet is to do all the ones in your book, then find some harder ones online. because in my experience. test questions are a lot harder than homework. they will have you factor stuff you have never seen before and make you think.

OpenStudy (anonymous):

thanks i think ill do that